OpenStudy (anonymous):

Ugh - please help me find the radius of this. x squared + y squared -4x-16y-9=0

OpenStudy (anonymous):

I know the center is (2,8)

OpenStudy (anonymous):

but finding the square root radius is giving me a headache

OpenStudy (anonymous):

r = √(9 + (16/2)² + (4/2)²) r = √(9 + 8² + 2²) r = √(9 + 64 + 4) r = √77